The time period of an earth satellite in circular & orbit is independent of
 
(a) both the mass and radius of the orbit
(b) ' radius of its orbit
(c) the mass of the satellite
(d) neither the mass of the satellite nor the radius of its orbit

Grade:6

1 Answers

Deepak Patra
9 years ago

(c) We have, mv2 / R+ x = GmM / ( R + x )2

x = height of satellite from earth surface

m = mass of satellite

? v2 = GM/( R+x ) or v = vGM/R+x

T = 2p (R + x) /v = 2p (R + x ) /vGM/R+x

This is independent of mass of satellite